OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

OpenStudy (nenadmatematika):

y-5=(3/2)(x-4) so y=(3/2)x-1

OpenStudy (anonymous):

THANK YOU

OpenStudy (nenadmatematika):

you're welcome :D